# Anisotropy can make a moving active fluid membrane rough or crumpled

### v1

@@ Abstract
We present a hydrodynamic theory of anisotropic and inversion-asymmetric moving active permeable fluid membranes. These are described by an anisotropic Kardar-Parisi-Zhang equation. Depending upon the anisotropy parameters, the membrane can be large-scale anisotropic and logarithmically rough with translational quasi long range order and orientational long range order, together with the relaxational dynamics being logarithmically faster than ordinary diffusion. For other choices of the anisotropy parameters, the membrane is either effectively isotropic and algebraically rough with translational short, but orientational long range order, or crumpled.
@@
The generality of the observations about the dynamical properties of the plasma membranes in live biological cells Alberts et al. (1994); Ramaswamy (2010); Joanny and Prost (2009); Marchetti et al. (2013); Levin and Korenstein (1991); Tuvia et al. (1997) has prompted scientists to construct minimal physical descriptions. It is, for instance, now understood that the macroscopic membrane properties in mixed lipid membranes depend on whether the membrane is in equilibrium Banerjee and Basu (2015), or active Banerjee and Basu (2018). In this Letter, we set up the hydrodynamic theory of a nearly flat patch of an anisotropic and inversion-asymmetric moving permeable active fluid membrane. The activity could be due to an active species living on the membrane, or due to activity in the embedding medium. For simplicity, we assume that there are no conservation laws or broken symmetry modes associated with the activity. Thus, if the activity is due to an active species, e.g., a membrane-spanning active protein, its concentration is not assumed to be conserved, or if it is due to the activity of the embedding medium, e.g., an orientable active fluid, it is in its isotropic phase. We consider an anisotropic membrane. For simplicity, we assume a fixed background for the membrane, i.e., the embedding fluid momentum is not conserved. We argue that the hydrodynamic equation for the undulation modes of such a membrane given by a single-valued height field $h({\bf x},t)$ is given by
$\frac{\partial h}{\partial t}+\frac{\lambda_{ij}}{2}\partial_{i}h\partial_{j}h=\nu_{ij}\partial_{i}\partial_{j}h+f.$ (1)
@@
Both $\lambda_{ij}$ and $\nu_{ij}$ are symmetric matrices. Noise $f$ has zero-mean and a variance given by
$\langle f({\bf x},t)f(0,0)\rangle=2D\delta^{2}(x)\delta(t).$ (2)
@@
In the absence of any detailed balance, the noise strength $D$ has no specific relation to the damping coefficient matrix $\nu_{ij}$. For arbitrary $\nu_{ij}$ and $\lambda_{ij}$, Eq. (11) is an anisotropic generalization of the well-known isotropic Kardar-Parisi-Zhang (KPZ) equation Barabási and Stanley (1995). For simplicity and to reduce the number of parameters, we further impose invariance of Eq. (11) separately under $(x,y)\rightarrow(\pm y,\pm x)$ in two dimensions (2D). We thus set $\nu_{xx}=\nu_{yy}=\nu>0$, $\nu_{xy}=\nu_{yx}=\overline{\nu}$, $\lambda_{xx}=\lambda_{yy}=\lambda$ and $\lambda_{xy}=\lambda_{yx}=\overline{\lambda}$, making it distinct from the one considered earlier Wolf (1991); Täuber and Frey (2002); Altman et al. (2015).
@@
The most prominent result from direct numerical simulations (DNS) of Eq. (11) in two dimensions (2D) is that it admits membrane conformations $h({\bf x},t)$ having strikingly different statistical properties controlled by the model parameters. See Figs. 1(a) and (b) and Figs. 2(a) and (b) for representative snapshots of $h({\bf x},t)$ falling in three distinct phases (color bars represent $h$ relative to its mean) depending on the degree of the surface anisotropy controlled by $\overline{\nu},\overline{\lambda}$; see also the corresponding movies MOV1 [for Fig. 1(a)], MOV2 [for Fig. 1(b)], MOV3(a) [for Fig. 2(a)] and MOV3(b)[for Fig. 2(b)] in Supplemental Material (SM) sup . We characterize these phases by numerically solving Eq. (11) and measuring width ${\mathcal{W}}\equiv\sqrt{\langle[h({\bf x},t)-\overline{h}(t)]^{2}\rangle}$, which as a function of time $t$ has an initial growing part, followed by a saturated part ${\cal W}_{\text{sat}}$ that fluctuates about a steady mean Barabási and Stanley (1995); $\overline{h}(t)$ is the mean height at time $t$. We find ${\mathcal{W}}_{\text{sat}}$ to scale as $(\ln L/a_{0})^{\tilde{\chi}}$ and $L^{\chi}$, respectively, for Fig. 1(a) and Fig. 1(b); $\tilde{\chi}\approx 0.43\pm 0.019,\chi=0.379\pm 0.008$ in our DNS studies, $L$ and $a_{0}$ are system size and small scale cutoff, respectively. See Fig. 3. Hence we name them logarithmic rough and algebraic rough phases respectively. For Fig. 2(a) and (b), no steady states are reached (within the computation time), giving ever increasing fluctuations, suggesting a complete loss of order, and hence we call it the crumpled phase. Snapshots in Figs. 1(c) and (d) for respectively the logarithmic and algebraic rough phases for different sets of model parameters over enhanced ranges as used in Figs. 2(a) and (b) for the crumpled phase indicate significantly larger fluctuations in the latter. See later for more details on DNS.
@@
Figure 1: Snapshots of the height profiles ($L=200$) in the steady states for (a) and (c) logarithmic rough phase, with parameters (a) $\nu=0.50$, $\overline{\nu}=0.25$, $\lambda=2$, and $\overline{\lambda}=4$, and (c) $\nu=0.60$, $\overline{\nu}=0.35$, $\lambda=5$, and $\overline{\lambda}=8$. Snapshots (b) and (d) correspond to the algebraic rough phase, with parameters (b) $\nu=0.50$, $\overline{\nu}=0.25$, $\lambda=24$, and $\overline{\lambda}=1$, and (d) $\nu=0.60$, $\overline{\nu}=0.35$, $\lambda=20$, and $\overline{\lambda}=2$. Notice the much larger fluctuations in the algebraic rough phase. See text.
Figure 2: Snapshots of $h({\bf x},t)$ ($L=200$) in the crumpled phase with $\nu=1,\lambda=2$ showing much larger fluctuations than those in Fig. 1. (a) $\overline{\nu}=2,\overline{\lambda}=2$ (with ripples at $\pi/4$), (b) $\overline{\nu}=-2,\overline{\lambda}=-2$ (with ripples at $43\pi/$ 4). See text.
We show that the logarithmic rough phase is characterized by translational quasi-long-range order (QLRO) and orientational long-range order (LRO) with relaxation time scales being logarithmically faster than ordinary diffusion. Such a surface is effectively inversion-symmetric but has large-scale anisotropy. Further, the algebraically rough phase has effective isotropy with inversion-asymmetry and translational SRO and orientational LRO, akin to a 2D KPZ surface Kardar et al. (1986); Tang et al. (1990); Barabási and Stanley (1995) and crumpled phase with translational and orientational SRO, respectively. These results should be contrasted with the corresponding isotropic case, which only has a rough phase.
We now derive Eq. (11) and use it to establish the above results and obtain the phase diagrams. The only hydrodynamic variable is $h({\bf x},t)$, measured with respect to an arbitrary base plane in the Monge gauge Nelson et al. (1989). We consider a nearly flat membrane patch to span at $z=0$ along the XY plane. Assuming simple relaxational dynamics, the general equation of motion of $h({\bf x},t)$, measured with respect to an arbitrary base plane in the Monge gauge Nelson et al. (1989), reads
$\frac{\partial h}{\partial t}=-\frac{\delta{\cal F}}{\delta h}+v_{\text{perm}}+f.$ (3)
Here, ${\cal F}$ is a free energy of an equilibrium fluid membrane, assumed tensionless, and $v_{\text{perm}}$ is the permeation velocity locally normal to the plane of the membrane. Here, ${\cal F}=\int d^{2}x[\frac{K}{2}(\nabla^{2}h)^{2}+C\nabla^{2}h]$, ignoring anharmonic contributions, is a free energy functional, which governs the relaxation of the membrane to equilibrium in the absence of any activity; here, $K>0$ is the bending stiffness and ${\bm{\nabla}}\equiv(\partial/\partial x,\,\partial/\partial y)$ is the 2D gradient operator; $C$ is the spontaneous curvature that can be generically present in an inversion-asymmetric membrane and makes the membrane curve one way or the other depending upon its sign. In the Monge gauge approximation, valid for a planar membrane patch, the spontaneous curvature term being a total derivative term vanishes and makes no contribution to $\cal F$.
@@
Consider a nearly flat patch of an inversion-asymmetric, anisotropic fluid membrane placed in an isotropic, active suspension of actin filaments Marchetti et al. (2013) grafted normally to the membrane locally. Let $\bf p$ be a three-dimensional unit vector ($p^{2}=1$) that gives the local orientation of the actin filaments and $\bf n$ be the local unit normal to the membrane; ${\bf n}=(\hat{z},-\nabla_{i}h)$ to the linear order in height fluctuations, in the present geometry with $i=x,y$. Locally normal grafting of the filaments implies ${\bf p}\cdot{\bf n}=1$. This yields $p_{j}=-\partial_{j}h\,(j=x,y)$ to the linear order in height fluctuations Sarkar and Basu (2013); Banerjee and Basu (2018). Now $v_{\text{perm}}$ Banerjee and Basu (2018), locally normal to an anisotropic membrane Barabási and Stanley (1995), reads
$v_{\text{perm}}=(v_{0}+\xi_{ij}\partial_{i}p_{j}+\tilde{\xi}_{ij}p_{i}p_{j})\sqrt{1+({\bm{\nabla}}h)^{2}}$ (4)
to the lowest order in small fluctuations $p_{i}$; $i=x,y$. In the cellular context, $v_{0}$ sets the scale of the drift speed of the membrane due to actin filament polymerization Sykes et al. (2010).The remaining terms in (4) give the drift due to local orientation fluctuations. Coefficient matrices $\xi_{ij}$ and $\tilde{\xi}_{ij}$ encode the anisotropy, which may originate from the anisotropy in the actomyosin cortex Mayer et al. (2010), or anisotropic inclusions in the membrane Fošnarič et al. (2005). Expanding and substituting for $p_{i}$ and retaining only the lowest order terms in gradients and nonlinearities, we obtain Eq. (11), where $\lambda_{ij}=\tilde{\xi}_{ij}$ is a measure of the average membrane speed due to fluctuations, and $\nu_{ij}=-\xi_{ij}$ is an anisotropic tension. Structures of these anisotropy parameters should depend explicitly on the specific model representation.
Figure 3: Log-log plots of (a) ${\mathcal{W}_{\text{sat}}}$ versus $\ln L$ ($a_{0}=1$) for $\alpha=0.50$, $\beta=2$ (logarithmic rough phase) and $L=100$, $300$, $500$, $800$, $1000$ and $1600$. Red dashed line represents linear fit to the numerical results displayed on a log-log scale with a slope $\tilde{\chi}=0.43\pm 0.019$. (b) ${\mathcal{W}_{\text{sat}}}$ with $L$ for $\alpha=0.50$, $\beta=0.042$ (algebraic rough phase) and $L=100$, $200$, $300$, $400$ and $500$. Red dashed line represents linear fit to the numerical results displayed on a log-log scale with a slope $\chi=0.379\pm 0.008$ (see text).
We now theoretically explain the above numerical observations, quantitatively characterizing the phases, by using Eq. (11). Equation (11) admits a pseudo-Galilean invariance under the transformation $t^{\prime}=t,\,x_{i}^{\prime}=x_{i}+\lambda_{ij}u_{j}t,\,h^{\prime}({\bf x}^{\prime},t^{\prime})=h({\bf x},t)+u_{i}x_{i}+\frac{\lambda_{ij}}{2}u_{i}u_{j}t$ sup . In (11), $\nu_{ij}$ plays the role of an (active) anisotropic surface tension. With $\nu>0$, linear stability requires that $\nu\pm\overline{\nu}\geq 0$, or $|\alpha|\leq 1$, i.e. $-1\leq\alpha\leq 1$; otherwise instability ensues. This means $\alpha_{c}=\pm 1$ is a threshold of $\alpha$, with larger $|\alpha|$ implies instability. This is the anisotropy-induced crumpling mentioned earlier in this Letter, which can occur even at zero noises. This instability threshold is independent of wavevector $q$, any ensuing crumpling actually takes place at all scales, meaning an arbitrarily small membrane will be crumpled, provided $|\alpha|>1$. This is reminiscent of the strongly crumpled phase in tensionless equilibrium asymmetric tethered membranes Banerjee et al. (2019a, b). In the crumpled phase, the variance of the local normal fluctuations $\sim\langle({\bm{\nabla}}h)^{2}\rangle$ diverges for a sufficiently large but finite $L$, giving translational and orientational SRO. In this phase, the growth rate is maximum at an angle $\pi/4$ and $43\pi/$ 4, respectively, with the $x$-axis for $\alpha>1$ and $\alpha<-1$. This is reflected in the ripples observed Fig. 2(a) and (b), and the corresponding movies MOV3(a) and MOV3(b) in SM sup .
We now turn to the uncrumpled case with $|\alpha|<1$, and consider its stability and fluctuation properties. In the linear theory with $|\alpha|<1$, width ${\mathcal{W}}_{\text{sat}}\sim\sqrt{\ln(L/a_{0})}$.As in the usual KPZ equation Kardar et al. (1986), the combination of noises and nonlinearities here can change the linear theory results substantially. To study the nonlinear effects in the present model (11), we perform a one-loop perturbative dynamic renormalization group (RG) analysis of Eq. (11) with $|\alpha|<1$. The Wilson dynamic RG method Täuber (2014); Forster et al. (1977) for our model closely resembles that for the KPZ equation Kardar et al. (1986); Barabási and Stanley (1995); Täuber (2014). Galilean invariance of (11) ensures that $\lambda_{ij}$ is unrenormalized, which is an exact statement and the analogue of the nonrenormalization of the nonlinear coupling constant in the usual KPZ equation Barabási and Stanley (1995); Frey and Täuber (1994). However, there are diverging one-loop corrections to $\nu,\,\overline{\nu}$ and $D$, which are represented by the one-loop Feynman diagrams (see SM) sup . Dimensional analysis allows us to identify an effective dimensionless coupling constant $g$ having a critical dimension 2 and a dimensionless ratio $\beta$ defined as
$g=\frac{\lambda^{2}D}{8\pi(\nu^{2}-\overline{\nu}^{2})^{\frac{3}{2}}},\;\beta=\frac{\overline{\lambda}}{\lambda},$ (5)
along with $\alpha$ defined earlier. The RG differential recursion relations ($\exp(l)$ is a dimensionless length) are
$\frac{dD}{dl}=D\bigg{[}z-2-2\chi+g\{\frac{(2+\alpha^{2})+\beta^{2}(1+2\alpha^{2})-6\alpha\beta}{2(1-\alpha^{2})}\}\bigg{]},$ (6)
$\frac{d\nu}{dl}=\nu[z-2-g\{\frac{\alpha\beta}{2}-\frac{\beta^{2}}{2}\}],$ (7)
$\frac{d\overline{\nu}}{dl}=\overline{\nu}[z-2-g\{\frac{1}{2}-\frac{\beta}{2\alpha}\}],$ (8)
$\frac{dg}{dl}=-g^{2}(\beta^{2}-1),$ (9)
together with $d\beta/dl=d/dl(\overline{\lambda}/\lambda)=0$. Here, $z$ and $\chi$ are the dynamic and roughness exponents, respectively Barabási and Stanley (1995)In the linear theory, i.e., with $\lambda_{ij}=0$, $z=2,\,\chi=0$, giving translational QLRO. Equation (9) gives that if $\beta^{2}>1$, then $g(l)\sim\frac{1}{(\beta^{2}-1)l}$ for large “RG time” $l$. On the other hand, if $\beta^{2}<1$, $g(l)$ diverges as $l\rightarrow l_{c}\sim{\cal O}(1)$ from below, where $l_{c}$ is a nonuniversal model parameter-dependent length. Of course, we cannot follow all the way to $g(l)$ diverging; as soon as $g(l)$ becomes ${\cal O}(1)$, the perturbation theory breaks down. This is the strong coupling algebraic rough phase or the model, whose scaling properties are not amenable to perturbative RG treatments, and is analogous to the strong coupling rough phase in the 2D KPZ equation Barabási and Stanley (1995); Tang et al. (1990).
The RG recursion relation of $\alpha$ reads
$\frac{d\alpha}{dl}=\frac{g}{2}\bigr{[}\alpha^{2}\beta-\alpha\beta^{2}-\alpha+\beta\bigr{]}.$
Setting $\frac{d\alpha}{dl}=0$, we obtain
$\alpha^{*}=1/\beta,\beta$
as the fixed points of $\alpha$. Thus in the perturbatively accessible region $\beta^{2}>1$, $\alpha^{*}=1/\beta$, $(\beta)$ is stable (unstable) fixed line. Focusing on the stable case $\beta^{2}>1$ and using the linear theory results $z=2$, $\chi=0$ we find $\nu(l)\sim\nu_{0}\sqrt{l},\,\overline{\nu}(l)\sim\overline{\nu}_{0}\sqrt{l},\,D(l)\sim D_{0}\sqrt{l}$ Mukherjee and Basu (2024), with $\nu_{0},\overline{\nu}_{0},D_{0}$ being the “initial” values of $\nu(l),\,\overline{\nu}(l),D(l)$ respectively. This gives logarithmically faster relaxation than ordinary diffusion, time-scale $t\sim r^{2}/\sqrt{\ln(r/a_{0})}$, where $r$ is a distance, meaning logarithmic breakdown of dynamic scaling. Similar logarithmically modulated noise strength and damping are also found in Refs. Haldar et al. (2023a, b); Jana et al. (2024). We further get
${\mathcal{W}}_{\text{sat}}\sim(D_{0}/{\mathcal{A}})\sqrt{\ln{L}/{a_{0}}},$ (10)
at the stable RG fixed point, where ${\mathcal{A}}\equiv[{4\pi\nu_{0}({1-(\frac{\overline{\nu}_{0}}{\nu_{0}})^{2}})^{1/2}}]^{1/2}$. This corresponds to a logarithmically rough surface with translational QLRO, statistically identical to a 2D Edward-Wilkinson (EW) surface Barabási and Stanley (1995); Edwards and Wilkinson (1982). An effective temperature $T_{\text{eff}}$ defined via $T_{\text{eff}}=D(l)/\nu(l)=D_{0}/\nu_{0}$, i.e., unchanged from its microscopic value. Nonetheless, due to its logarithmically modulated noise and damping, at this RG fixed point our model is different from the 2D EW equation. While the equal-time height-difference correlator is same as in the EW equation, the more general time-displaced correlator is different. With $\alpha^{*}=1/\beta$ and $\beta^{2}>1$ at the RG fixed point, the logarithmic rough phase is macroscopically anisotropic.
In the perturbatively inaccessible case with $\beta^{2}<1$, we have $\alpha^{*}=\beta$ as the stable fixed point, implying large-scale isotropy. In this case $d\nu/dl,d\overline{\nu}/dl=0,dD/dl>0$, reminiscent of the RG flow equations in the 2D isotropic KPZ equation Barabási and Stanley (1995).

### v2

@@ Abstract
We present a hydrodynamic theory of anisotropic and inversion-asymmetric moving active permeable fluid membranes. These are described by an anisotropic Kardar-Parisi-Zhang equation. Depending upon the anisotropy parameters, the membrane is either effectively isotropic and algebraically rough with translational short, but orientational long range order, or unstable, suggestive of membrane crumpling.
@@
The generality of the observations about the dynamical properties of the plasma membranes in live biological cells Alberts et al. (1994); Ramaswamy (2010); Joanny and Prost (2009); Marchetti et al. (2013); Levin and Korenstein (1991); Tuvia et al. (1997) has prompted scientists to construct minimal physical descriptions. It is, for instance, now understood that the macroscopic membrane properties in mixed lipid membranes depend on whether the membrane is in equilibrium Banerjee and Basu (2015), or active Banerjee and Basu (2018). In this Letter, we set up the hydrodynamic theory of a nearly flat patch of an anisotropic and inversion-asymmetric moving permeable active fluid membrane. The activity could be due to an active species living on the membrane, or due to activity in the embedding medium. For simplicity, we assume that there are no conservation laws or broken symmetry modes associated with the activity. Thus, if the activity is due to an active species, e.g., a membrane-spanning active protein, its concentration is not assumed to be conserved, or if it is due to the activity of the embedding medium, e.g., an orientable active fluid, it is in its isotropic phase. We consider an anisotropic membrane. We argue that the, assuming Rouse dynamics Cai and Lubensky (1995) for simplicity, hydrodynamic equation for the undulation modes of an anisotropic membrane given by a single-valued height field $h({\bf x},t)$ is given by
$\frac{\partial h}{\partial t}+\frac{\lambda}{2}{({\bm{\nabla}}h)^{2}}=\nu_{ij}\partial_{i}\partial_{j}h+f.$ (1)
@@
Here $\nu_{ij}$ is symmetric under the interchange of $i$ and $j$, and is an effective anisotropic tension. Noise $f$ has zero-mean and a variance given by
$\langle f({\bf x},t)f(0,0)\rangle=2D\delta^{2}({\bf x})\delta(t).$ (2)
@@
In the absence of any detailed balance, the noise strength $D$ has no specific relation to the damping coefficient matrix $\nu_{ij}$. For arbitrary $\nu_{ij}$, Eq. (1) is an anisotropic generalization of the well-known isotropic Kardar-Parisi-Zhang (KPZ) equation Barabási and Stanley (1995). Since in general $\nu_{ij}\neq 0$ for $i\neq j$, Eq. (1) is anisotropic. We set $\nu_{xx}=\nu_{x}$, $\nu_{yy}=\nu_{y}$, with $\nu_{x}\neq\nu_{y}$ both positive and $\nu_{xy}=\nu_{yx}=\overline{\nu}$ can be of any sign. Equation (1) is invariant under the joint transformation $(x,y)\to(-x,-y)$, which is of lower symmetry than the one considered earlier Wolf (1991); Täuber and Frey (2002); Altman et al. (2015). See also Ref. da Silveira and Kardar (2003) for detailed studies on the stationary states of anisitropic KPZ equations.
@@
The most prominent result from direct numerical solutions (DNS) of Eq. (1) in 2D is that it admits different membrane conformations $h({\bf x},t)$ having strikingly distinct statistical properties controlled by the model parameters. See Fig. 1 for representative snapshots of $h({\bf x},t)$ falling in two distinct phases (color bars represent $h$ relative to its mean) depending on the degree of the surface anisotropy controlled by $\nu_{x},\nu_{y},\overline{\nu}$. The corresponding movies MOV1 [for Fig. 1(a)], MOV2 [for Fig. 1(c)] and MOV3 [for Fig. 1(d)] are available in Supplemental Material (SM) sup . We characterize these phases by numerically solving Eq. (1) and measuring width ${\mathcal{W}}\equiv\sqrt{\langle[h({\bf x},t)-\overline{h}(t)]^{2}\rangle}$, by averaging over different realisations, which as a function of time $t$ has an initial growing part, followed by a saturated part ${\cal W}_{\text{sat}}$ that fluctuates about a steady mean Barabási and Stanley (1995); $\overline{h}(t)$ is the spatially averaged mean height at time $t$. We find ${\mathcal{W}}_{\text{sat}}$ to scale as $L^{\chi}$, respectively, for Fig. 1(a); $\chi=0.381\pm 0.002$; in our DNS studies, $L$ is system size. See Fig. 3. Hence we name it algebraic rough phase. For Fig. 1(c) and (d), no steady states are reached (within the computation time), giving ever increasing fluctuations, suggesting instability and a complete loss of order. Snapshot in Fig. 1(b) represent the algebraic rough phase for a different set of model parameters over enhanced ranges as used in Figs. 1(c) and (d) for the crumpled phase, indicate significantly larger fluctuations in the latter. See also later for the detailed numerical results from our DNS studies.
@@
Figure 1: Snapshots of the height profiles ($L=200$) in the steady states for (a) and (b) algebraic rough phase, with parameters (a) $\nu_{x}=0.50$, $\nu_{y}=0.45$, $\overline{\nu}=0.10$, $D=0.005$ and $\lambda=24$, and (b) $\nu_{x}=0.50$, $\nu_{y}=0.55$, $\overline{\nu}=0.20$, $D=0.005$ and $\lambda=24$. Snapshots (c) and (d) correspond to the crumpled phase, with parameters (c) $\nu_{x}=1.0$, $\nu_{y}=0.20$, $\overline{\nu}=1.0$, $\nu_{4}=0.10$, $D=1.0$ and $\lambda=2.20$, (with ripples at an arbitrary angle) and (d) $\nu_{x}=1.0$, $\nu_{y}=1.0$, $\overline{\nu}=2.0$, $\nu_{4}=0.10$, $D=1.0$ and $\lambda=2.30$ (with ripples at $\pi/4$). Notice the much smaller fluctuations in the algebraic rough phase, compared to the crumpled phase. (See text).
We show that our model exhibits algebraic rough phase and is characterized by translational short-range order (SRO) and orientational long-range order (LRO) and specifically belongs to the KPZ universility class. Using DNS we verify for different sets of model parameters that the scaling exponents remain almost same indicating the putative algebraic rough phase is model parameter independent and indeed belongs to the KPZ universility class. Further, the algebraically rough phase is effectively isotropic with inversion-asymmetry and translational short range order (SRO) and orientational LRO, akin to a 2D KPZ surface Kardar et al. (1986); Tang et al. (1990); Barabási and Stanley (1995) and the crumpled phase has translational and orientational SRO, respectively.
We now derive Eq. (1) and use it to establish the above results and obtain the phase diagrams. The membrane configurations are given by three-dimensional position vector ${\bf R}({\vec{s}},t)$, where ${\vec{s}}=(s_{1},s_{2})$ is a two-dimensional position vector embedded on the membrane, which label points on the membrane. The local membrane velocity ${\bf v}({\vec{s}},t)\equiv\partial{\bf R}/\partial t$, for a convected or embedded coordinate system Aris (1989); Hu et al. (2007). We now specialize for nearly flat inversion-asymmetric active membrane patches without any overhangs, for which the Monge gauge Nelson et al. (1989) is a convenient choice, in which $(s_{1},s_{2})\equiv(x,y)$ and ${\bf R}({\bf x})=({\bf x},h({\bf x})),\,{\bf x}=(x,y)$. The only hydrodynamic variable is $h({\bf x},t)$, measured with respect to an arbitrary base plane in the Monge gauge Nelson et al. (1989). Due to the breakdown of inversion symmetry, the patch, considered permeable, is generically moving in a direction local normal to itself. A change in the membrane conformation $h({\bf x},t)$ can take place by the drift of the membrane due to activity and relaxational dynamics, which we assume to be Rouse dynamics for simplicity. The general equation of motion of $h({\bf x},t)$, measured with respect to an arbitrary base plane in the Monge gauge Nelson et al. (1989), reads
$\frac{\partial h}{\partial t}=v\sqrt{1+({\bm{\nabla}}h)^{2}}-\Gamma\frac{\delta F}{\delta h}+f,$ (3)
[…]
$v(h)\equiv v_{0}+\nu_{ij}\partial_{i}\partial_{j}h.$ (4)
In the cellular context, $v_{0}$ sets the scale of the drift speed of the membrane, e.g., due to actin filament polymerization Sykes et al. (2010) (see also below). The second term on the rhs of (4) gives the contribution from the membrane conformation fluctuations to the drift. Assuming small fluctuations, we expand the square root in (3) and obtain (1).
@@
A possible microscopic realization of this would be a nearly flat patch of an inversion-asymmetric, anisotropic fluid membrane placed in an isotropic, active suspension of actin filaments Marchetti et al. (2013) grafted normally to the membrane locally. Let $\bf p$ be a three-dimensional unit vector ($p^{2}=1$) that gives the local orientation of the actin filaments and $\bf n$ be the local unit normal to the membrane; ${\bf n}=(\hat{z},-\nabla_{i}h)$ to the linear order in height fluctuations ($i=x,y$). Locally normal grafting of the filaments implies ${\bf p}\cdot{\bf n}=1$, giving $p_{i}=-\partial_{i}h$ ($i=x,y$) and $p_{z}=1$ to the linear order in fluctuations. The actin filaments exert forces locally normally to the membrane Chen and Mikhailov (2010), giving a three-dimensional local permeation flow velocity
${\bf v}\sim v_{0}{\bf p}+{\bm{\nabla}}\cdot[{\bm{\zeta}}{\bf p}{\bf p}].$ (5)
Now, extracting the component of $\bf v$} along the membrane’s local normal and to linear order in $h$ yields $v$ as above Maitra et al. (2014); Banerjee and Basu (2018). Writing the second term on the rhs of (5) as the divergence of an active stress ${\bm{\sigma}}^{a}\equiv{\bm{\zeta}}{\bf p}{\bf p}$ i.e. $\sigma_{ij}^{a}\equiv\zeta_{ij\alpha\beta}p_{\alpha}p_{\beta}$, we see that $\nu_{ij}$ is proportional to $\sigma_{ij}^{a}$. Matrix ${\bm{\zeta}}\equiv\zeta_{ij\alpha\beta}$ encodes the anisotropy in the membrane, e.g., anisotropy in the actomyosin cortex Mayer et al. (2010), or anisotropic inclusions in the membrane Fošnarič et al. (2005). The forms of these anisotropy parameters should depend explicitly on the specific model representation. In the absence of anisotropy, in which case the strength of the active stress is described by a scalar parameter, say $\Delta\mu$, a positive (negative) $\Delta\mu$ refers to extensile (contractile) activity Marchetti et al. (2013). Here, however the active stress is characterized by three parameters, and hence a simple extensile/contractile characterization is no longer possible.